Define the type of given PDE and find the solution. (y^2+z^2-x^2 )p+2xyq+2zx=0


1
Expert's answer
2020-10-15T19:56:14-0400

Given differential equation is "(y^2+z^2-x^2 )p+2xyq+2zx=0"


Then we can write,

"\frac{dx}{(y^2+z^2-x^2 )}=\frac{dy}{2xy}=\frac{dz}{-2zx}"


Taking last two terms,

"\frac{dy}{2xy}=\frac{dz}{-2zx}"


Integrating both sides, we get

"\int\frac{dy}{y}=\int \frac{dz}{-z}"

"log |y| = -log|z| + logC_1 \implies yz = C_1" (1)

then "z = \frac{C_1}{y}"


taking first two terms,

"\frac{dx}{(y^2+z^2-x^2 )}=\frac{dy}{2xy}"


putting value of z from equation (1),

"\frac{dx}{(y^2+\frac{C_1^2}{y^2}-x^2 )}=\frac{dy}{2xy}"


"{(y^2+\frac{C_1^2}{y^2}-x^2 )}{dy}={2xy}{dx}"

"{2xy}{dx} - {(y^2+\frac{C_1^2}{y^2}-x^2 )}{dy} = 0" (2)

This equation is exact differential equation, as

"\frac{\partial M}{\partial y} = \frac{\partial N}{\partial x} = 2x"

where "M ={2xy}, N = - {(y^2+\frac{C_1^2}{y^2}-x^2 )}"

then integrating both sides equation (2),

"\int{2xy}{dx} - \int{(y^2+\frac{C_1^2}{y^2}-x^2 )}{dy} = C_2"


"x^2y-\frac{y^3}{3}+\frac{C_1^2}{y} = C_2"

putting value of "C_1"

"x^2y-\frac{y^3}{3}+yz^2 = C_2" (3)



Hence equation (1) and (3) are the required solutions of the equation.